# SpiralShard: Highly Concurrent and Secure Blockchain Sharding via Linked Cross-shard Endorsement

### v1

@@ Abstract
To enhance the concurrency of blockchain sharding securely, we propose SpiralShard. The intuition is to allow the existence of some shards with a larger fraction of malicious nodes (i.e., corrupted shards), thus reducing shard sizes. SpiralShard can configure more and smaller shards for higher concurrency at the same network size. To ensure security with the existence of corrupted shards, we propose the Linked Cross-shard Endorsement (LCE) protocol. According to our LCE protocol, the blocks of each shard are sequentially verified and endorsed by a group of shards before being finalized. As a result, a corrupted shard can eliminate forks with the help of the other shards. We implement SpiralShard based on Harmony and conduct extensive evaluations. Experimental results show that, compared with Harmony, SpiralShard achieves around 19$\times$ throughput gain under a large network size with 4,000+ nodes.
@@ I Introduction
Blockchain has emerged as a critical technology to support distributed applications [2, 4]. However, traditional blockchain systems [2, 4] fall short regarding scalability. Sharding is a promising approach to improving blockchain scalability [29, 46]. It partitions nodes in the blockchain network into smaller committees (shards). Each shard maintains a subset of the whole blockchain ledger state and conducts intra-shard consensus to generate blocks in parallel. Therefore, systems can set more shards with sufficient nodes to process transactions, improving scalability.
However, to preserve security, existing permissionless blockchain sharding systems usually configure large size for each shard, causing poor transaction concurrency [6, 21, 23, 29, 35, 46, 47]. In general, under the same network size, we can increase the number of shards by reducing the shard size (i.e., the number of nodes per shard) to improve the transaction concurrency [8]. However, arbitrarily reducing shard size will cause security problems. Specifically, most existing blockchain sharding systems require that the fraction of malicious nodes in each shard be less than $1/3$ (i.e., the fault tolerance threshold of BFT-typed consensus [12]). As nodes are typically randomly assigned to shards, merely reducing shard size can easily result in a larger fraction of malicious nodes in certain shards. In this case, these shards are corrupted, and the malicious nodes can launch attacks (e.g., double-spending [2]) to compromise security. Therefore, most sharding systems have to set large shard size to ensure a negligible failure probability for each shard. In OmniLedger [29], a shard must include $600$ nodes when 1/4 of the network nodes are malicious. This large shard size undermines intra-shard consensus efficiency and diminishes the shard count, consequently limiting transaction concurrency.
@@
In this paper, we aim to boost transaction concurrency for blockchain sharding securely and efficiently. This is achieved by proposing a blockchain sharding system named SpiralShard. Unlike most previous works, which require each shard to be uncorrupted, our intuition is to allow the existence of some corrupted shards with a larger fraction of malicious nodes. Therefore, the system has more and smaller shards for higher concurrency. To prevent corrupted shards from compromising the system’s security, we design an Linked Cross-shard Endorsement (LCE) protocol. Specifically, shards are divided into multiple endorsement groups, each belonging to one group. Each shard must include block headers from other shards within its blocks to express its endorsement decision. Blocks within a shard require endorsement in sequence from all other shards within the same endorsement group to be finalized securely.
With the help of endorsement groups, each shard can securely finalize its blocks, even if there are forks. As a result, the shard size is reduced without violating security, and the concurrency is boosted.
However, it is not straightforward to design the LCE protocol. We summarize the main challenges as follows.
@@
1). How to maintain low overheads for each shard during cross-shard verification? Our LCE protocol leverages cross-shard verification and endorsement to secure corrupted shards, introducing additional overhead. For example, to validate blocks from other shards, a straightforward idea is to store the state of other shards and verify raw transactions, leading to higher storage, computation, and communication overhead for each node. Hence, we need to reduce the overhead of cross-shard verification and endorsement.
To overcome this considerable overhead, our main idea is that each shard verifies and ensures its transactions’ validity by using its own shard’s state (rather than storing the state of others). On the other hand, the LCE protocol addresses the possible forking problem by transmitting block headers across the shards, reducing overhead. To achieve this goal, we first observed that when the number of malicious nodes is fewer than the quorum size (representing the votes needed for BFT-typed consensus, set as $2/3$ of the shard size), blocks with invalid transactions fail to amass enough votes during intra-shard consensus. Hence, the validity of each block can be verified by checking the votes in the block header, and malicious nodes cannot manipulate transactions to compromise safety. Therefore, we limit the fraction of malicious nodes in corrupted shards to be less than $2/3$, guaranteeing the transaction’s validity. In this case, each shard only needs to transmit its block headers across shards, and another shard can address the forking problem and verify the block’s validity via the block header but not the raw transactions. This mechanism significantly reduces the overhead for each node.
@@
2). How to preserve safety and liveness in LCE protocol while some shards are corrupted? In SpiralShard, the safety indicates that honest nodes agree on the same and valid shard chain for each shard. Although the validity of blocks can be verified by checking the votes in the block header, a corrupted shard may fork its chain to attack safety. Considering that, our intuition is to leverage an honest shard to endorse the blocks of corrupted shards. Since such an honest shard will not have forks, it will only endorse blocks on one fork branch in the corrupted shard chain, thus eliminating the forks. However, we cannot know in advance whether a shard is honest since nodes are randomly assigned to shards. Therefore, we establish the concept of the endorsement group, which comprises multiple shards (while each shard simultaneously belongs to an individual endorsement group). Additionally, we fine-tune the size of the endorsement group to guarantee that each group has at least one honest shard. Subsequently, we require each shard of an endorsement group to endorse the blocks of other shards. In this case, once a corrupted shard receives endorsements from all shards in its endorsement group, only one fork branch is selected, eliminating forks and ensuring safety.
However, in the aforementioned design, the asynchronous endorsements of the same corrupted shard by different shards within the same endorsement group give rise to a particular challenge. This setup may result in a scenario where various honest shards endorse distinct blocks of the identical corrupted shard, thereby preventing any block from accumulating sufficient endorsements to be finalized. This situation threatens the system’s liveness. We require the shards in an endorsement group to conduct endorsement sequentially to tackle this issue. Each shard exclusively endorses the preceding shard in the endorsement group, following the ascending order of the shard IDs. Simultaneously, the endorsement from a shard implies its support for the endorsement decision made by the preceding shard. Consequently, endorsements from different shards are linked, contributing to the finality of the same original block. In other words, subsequent honest shards can only endorse the endorsements made by preceding honest ones and cannot contradict them. This mechanism guarantees the safe finalization of a corrupted shard’s block once all shards in an endorsement group have sequentially declared their endorsements.
@@
To further refine our liveness guarantee against the possible liveness attack by corrupted shards (i.e., not generating any blocks and halting the linked endorsement), we adopt a mixed threat model where Byzantine nodes target both safety and liveness while alive-but-corrupt nodes [36] (CCS’ 19) only compromise safety but not liveness. Under this mixed threat model, we ensure that the fraction of Byzantine nodes in each shard remains below $1/3$, and we allow corrupted shards to contain additional no more than $1/3$ of a-b-c nodes. These measures are guided by our rigorous theoretical analysis presented in Section V. As a result, the quorum size of votes can be achieved to advance consensus within each shard, ensuring the continuous extension of the linked endorsement.
3). How to maintain high efficiency when there are some slow shards in the system? In SpiralShard, we pipeline the block generation to enhance overall throughput. Generally, shards may have different block generation speeds, as the network is usually not synchronized. Meanwhile, the sequentially processed cross-shard endorsement may be halted due to slow shards. To solve this problem, we design a pipelining mechanism to allow each shard to generate new blocks optimistically. Therefore, each shard can generate new blocks before its blocks are finalized. To ensure the security of the proposed pipelining mechanism, we require that the generated blocks be finalized after being endorsed by the whole endorsement group. To preserve efficiency, our pipelining mechanism requires a shard to endorse the latest block instead of every block of the preceding shard. With this design, when a block is endorsed by all shards from its group, this block and its parent blocks are all finalized.
With the above challenges addressed, this paper mainly has the following contributions.
@@
  • We implement SpiralShard based on Harmony [5] and conduct evaluations based on large-scale Amazon EC2 deployment. Compared with Harmony, SpiralShard achieves around 19$\times$ throughput under a large network size with 4,000+ nodes.
@@ II-A Blockchain and Blockchain Sharding
Traditional blockchain systems require nodes to store and verify every transaction on a single chain, resulting in poor performance and scalability. For example, Bitcoin can only process 7 tx/s (transaction per second), while Ethereum can only process 15 tx/s [24]. Recent works adopt sharding [35] to improve the scalability of blockchain [35, 29, 46, 6, 17, 45, 31, 38, 25]. The main idea of sharding is to partition the nodes into multiple small committees (shards). Each shard maintains a disjoint subset of the whole blockchain state, conducts intra-shard consensus, and processes distinct transactions in parallel. However, existing blockchain sharding systems still suffer from limited concurrency.
@@ II-B Shard Size and Concurrency
Some works [35, 29, 46, 17, 21, 23, 47] try to improve system concurrency by reducing the shard size. However, they have various limitations. For instance, some systems [35, 29, 46, 23, 47] tolerate a larger fraction (i.e., $<1/2$) of malicious nodes within each shard based on a less practical assumption that the network within each shard is synchronous. However, real-world network environments are often not synchronous (e.g., partial-synchronous), especially for large-scale blockchain systems. SGX-Sharding [17] reduces the shard size by applying trusted hardware. However, it brings additional deployment overhead to each node, limiting its generality. Some other solutions reduce the shard size at the cost of reducing the systems’ overall resiliency, compromising the system’s security. For example, in Pyramid [21], the assumption entails a smaller fraction of malicious nodes in the network, resulting in a system capable of withstanding less than $1/8$ of malicious nodes.
@@ II-C Blockchain Sharding with Corrupted Shards
Only a handful of blockchain sharding studies allow corrupted shards, with various limitations. Free2Shard [39] allows corrupted shards and preserves the system’s security via a network-wide consensus. However, it is based on the assumption of a global synchronous network environment. Therefore, it does not work correctly under a practical partial-synchronous network. In addition, Free2Shard adopts the Proof of Work (PoW) consensus protocol, which does not provide deterministic finality and is prone to forking problems. In GearBox [18], the system relies on a network-wide consensus to ensure the system’s security, which can easily lead to a performance bottleneck in large-scale networks. The latest work, known as CoChain [32], similarly stems from an intuition that allows shard corruption to enhance concurrency. In CoChain, shards are randomly assigned to CoC (Consensus on Consensus) groups. However, CoChain only permits the failure of less than one-third of the shards within a group. When the group size is no larger than 3, CoChain cannot allow any corrupted shard, losing feasibility. More importantly, CoChain needs to rely on very time-consuming state migration to replace the corrupted shards when they fork. In contrast, our SpiralShard can utilize the LCE protocol to easily help corrupted shards remove excess forks without the need for shard replacement or recovery.
@@
SpiralShard is a novel design that is different from existing blockchain sharding systems. Built upon the assumption of partial-synchronous networks, it is more practical. Leveraging BFT-typed intra-shard consensus ensures finality and energy efficiency. Unlike systems reliant on global consensus to resolve forks, it leverages multiple endorsement groups, preserving scalability. It requires at least one shard of each group to be honest, offering greater flexibility in endorsement group size.
II-D Interchain Timestamping
In the cross-blockchain context, timestamping technology enables less secure blockchains to harness the security of more robust counterparts. Specifically, producer chains are responsible for recording block headers of consumer chains. As long as these producer chains remain secure and un-forked, the sequence of blocks in the consumer chain remains consistent, ensuring security. Tas et al. [44] introduce an interchain timestamping approach involving multiple producer chains recording blocks from a consumer chain. This allows the consumer chain to leverage security assurances from all the producer chains. While Tas’s proposal shares conceptual similarities with the LCE protocol in SpiralShard, it is not tailored for a sharding-based blockchain scenario. Moreover, it lacks a rigorous computation of system failure probabilities and cannot guarantee liveness in some cases.
III System and Threat Model
@@ III-A Network Model
SpiralShard works on a partial-synchronous Peer to Peer (P2P) network, in which each message will reach the receiver within an unknown upper bound $\delta$ delay after being sent from nodes [20]. Like most previous blockchain sharding systems, the messages are propagated through gossip protocol [19]. Each node in the system has its public/secret key pair, which can be authenticated through a trusted Public-Key Infrastructure (PKI) to represent its identity while sending messages. Like many prior works [21, 29, 5], SpiralShard utilize a beacon chain to publicly store the identities of participants.
@@
In SpiralShard, the network consists of $N$ nodes. The shard size (i.e., the number of nodes per shard) and the endorsement group size (i.e., the number of shards per endorsement group) are denoted as $S$ and $G$, respectively. This implies that each endorsement group contains $S\cdot G$ nodes, with $N/S$ shards within the network. A node’s endorsement group identity is determined by its shard identity (i.e., $GroupID=\lfloor ShardID/G\rfloor$).
III-B Transaction Model
@@
SpiralShard employs the account/balance transaction model, widely adopted by existing works to represent ledger states [21, 5, 45, 11, 33, 17, 7, 26, 22]. In SpiralShard, a specific account’s state (e.g., balance) is maintained by one shard, determined by the hash of its account address. Accordingly, a transaction in the network is routed to the corresponding shard based on its associated account address. It is important to emphasize that this includes both intra-shard and cross-shard transactions. Besides, the two accounts involved in cross-shard transactions may be in different endorsement groups.
III-C Threat Model

### v2

@@ Abstract
[…]
@@ I Introduction
Blockchain has emerged as a critical technology to support distributed applications [bitcoin, eth]. However, traditional blockchain systems [bitcoin, eth] fall short regarding scalability. Sharding is a promising approach to improving blockchain scalability [omniledger, rapidchain, li2025sp]. It partitions nodes in the blockchain network into multiple disjoint committees (shards). Each shard maintains a subset of the whole blockchain ledger state and conducts intra-shard consensus to generate blocks and process transactions in parallel. In general, under the same network scale, configuring a large amount of small shards helps to achieve higher transaction concurrency and throughput (i.e., transactions per second, TPS) [divideAndScale, cochain, gearbox].
However, naively configuring numerous small shards can easily lead to certain shards becoming corrupted, compromising security. To ensure security, existing permissionless blockchain sharding systems typically adopt relatively large shard sizes [chainspace, pyramid, omniledger, elastico, rapidchain, cycledger, li2025sp]. Under the same network scale, such larger shard sizes significantly reduce the total number of shards in the network and diminish the efficiency of intra-shard consensus, consequently severely restricting the transaction concurrency (TPS) of sharding-based blockchain systems. Specifically, nodes within blockchain sharding systems (especially permissionless ones) are typically allocated to shards randomly [elastico, rapidchain, monoxide, li2022jenga, LBCHAIN, li2025sp]. Due to this randomness, when the network size and overall malicious node ratio remain unchanged, naively adopting small shard sizes can easily lead to excessively high malicious node ratios in some shards, causing shard corruption (e.g., $\geq 1/3$ malicious nodes for BFT-type (Byzantine Fault Tolerance) consensus under partial-synchronous network [castro1999practical]) and compromising system security [bitcoin]. Therefore, most sharding systems have to set large shard size to ensure a negligible failure probability for each shard, guaranteeing security. For example, in OmniLedger [omniledger], a shard could include $600$ nodes when 1/4 of the network nodes are malicious.
@@
In this paper, we aim to boost transaction concurrency for blockchain sharding securely and efficiently. This is achieved by proposing a blockchain sharding system named SpiralShard. Unlike most previous works, which require each shard to be uncorrupted, our intuition is to allow the existence of some corrupted shards with a larger fraction of malicious nodes. Therefore, the system has more and smaller shards for higher concurrency. To prevent corrupted shards from compromising system security, we design an Linked Cross-shard Endorsement (LCE) protocol. Specifically, shards are divided into multiple endorsement groups, each belonging to one group. Each shard must include block headers from other shards within its blocks and reach intra-shard consensus to express its endorsement decision. Blocks within a shard require endorsement in sequence from all other shards within the same endorsement group to be finalized securely. With the help of endorsement groups, each shard can securely finalize its blocks, even if there are forks. As a result, the shard size is reduced without violating security, and the concurrency is boosted.
It is not straightforward to design the LCE protocol. We summarize the main challenges as follows.
@@
1). How to maintain low overheads for each shard during cross-shard verification? The LCE protocol leverages cross-shard verification and endorsement to secure corrupted shards, introducing additional overhead. To validate blocks from other shards, a straightforward idea is to store the state of other shards and verify raw transactions, leading to higher storage, computation, and communication overhead for each node. Hence, we need to reduce the overhead of cross-shard verification and endorsement.
To reduce overhead, our core idea is that each shard internally ensures the validity of individual transactions through consensus (preventing transaction manipulation), and subsequently employs the LCE protocol to address potential fork problems. Consequently, cross-shard verification of individual transactions is unnecessary; instead, shards only need to exchange lightweight block headers to detect forks. To realize this goal, we rigorously conduct probabilistic modeling and derive explicit mathematical formulations (in Section VI-A). Based on this modeling, we precisely fine-tune shard sizes to ensure—except with negligible failure probability (formally defined in Section IV-B)—that the proportion of malicious nodes within each shard is always below 2/3. This threshold is selected because our analysis shows that when the fraction of malicious nodes is below the quorum size (i.e., 2/3 for BFT consensus protocols in partially synchronous networks) required for consensus, invalid transactions cannot accumulate enough malicious signatures to pass the consensus procedure. Hence, under such conditions, each shard inherently guarantees transaction validity via internal consensus. However, corrupted shards may still produce forks that cannot be resolved independently. In this scenario, the LCE protocol efficiently resolves forks solely by monitoring block headers, thereby maintaining low overhead.
@@
2). How to preserve safety and liveness in LCE protocol while some shards are corrupted? As discussed above, some randomly corrupted shards may threaten system safety (i.e., honest nodes agree on the same and valid shard chain for each shard) by creating forks. To securely address the fork issue, we first require multiple shards to form an endorsement group, within which shards perform mutual verification and endorsement. More importantly, through rigorous probabilistic analysis, we carefully adjust the size of each endorsement group, enabling the system to include as many endorsement groups as possible while simultaneously ensuring (except with negligible failure probability) that each endorsement group contains at least one honest shard. Consequently, honest shards within an endorsement group guarantee the honest selection of only one single fork for endorsement. This mechanism ensures that no shard can have multiple conflicting forks being finalized, as no shard can have multiple forks endorsed by all members of its endorsement group.
However, the aforementioned design could result in different honest shards within an endorsement group endorsing distinct forks. Consequently, no single fork would obtain endorsements from all group members, preventing any fork from being finalized and thereby impairing system liveness (the system does not stall indefinitely). To address this issue, we enforce a predetermined endorsement order (e.g., ascending order of the shard IDs) among the shards within each endorsement group, whereby each shard endorses only the decision previously made by its predecessor shard. Consequently, the endorsement from each shard implicitly validates the endorsements provided by all preceding shards. Under this mechanism, each shard’s consensus outcome—reflected in the block header—accumulates a sequential chain of linked endorsements from the endorsement group members. This sequentially linked endorsement strategy ensures that honest shards invariably endorse the same fork consecutively, enabling a single fork to receive unanimous endorsements from all endorsement group members and thereby achieve finality.
@@
To further enhance system liveness against possible liveness attacks by corrupted shards (i.e., hindering the advancement of the intra-shard consensus or LCE protocol), we adopt a mixed threat model comprising both Byzantine nodes, which simultaneously compromise safety and liveness, and alive-but-corrupt (a-b-c) nodes [flexible] (CCS ’19), which maintain liveness but exclusively attack safety. Under this mixed threat model, we ensure that the fraction of Byzantine nodes in each shard remains below $1/3$, and we allow corrupted shards to contain additional no more than $1/3$ of a-b-c nodes. These measures are also guided by our rigorous theoretical analysis and shard size fine-tuning. In this case, the continued advancement of the intra-shard consensus and LCE protocol is guaranteed, refining liveness.
3). How to maintain high efficiency when there are some slow shards in the system? In SpiralShard, we pipeline the block generation to enhance overall throughput. Generally, shards may have different block generation speeds. Meanwhile, the sequentially processed cross-shard endorsement may be halted due to slow shards. To solve this problem, we design a pipelining mechanism to allow each shard to generate new blocks optimistically. Therefore, each shard can generate new blocks before its blocks are finalized. To ensure the security of the proposed pipelining mechanism, we require that the generated blocks be finalized only after being endorsed by the whole endorsement group. To preserve efficiency, our pipelining mechanism requires a shard to endorse the latest block instead of every block of the preceding shard. With this design, when a block is endorsed by all shards from its group, this block and its parent blocks are all finalized.
This paper mainly has the following contributions:
@@
[…]
@@ II-A Blockchain and Blockchain Sharding
Traditional blockchain systems require nodes to store and verify every transaction on a single chain, resulting in poor performance and scalability. For example, Bitcoin can only process 7 tx/s (transaction per second), while Ethereum can only process 15 tx/s [huang2021survey]. Recent works adopt sharding [elastico] to improve the scalability of blockchain [elastico, omniledger, rapidchain, chainspace, sgxSharding, monoxide, li2022jenga, LightCross, accountMigra, li2025sp]. The main idea of sharding is to partition the nodes into multiple small committees (shards). Each shard maintains a disjoint subset of the whole blockchain state, conducts intra-shard consensus, and processes distinct transactions in parallel. However, existing blockchain sharding systems still suffer from limited concurrency.
@@ II-B Shard Size and Concurrency
Some works [elastico, omniledger, rapidchain, sgxSharding, pyramid, cycledger] try to improve system concurrency by reducing the shard size. However, they have various limitations. For instance, some systems [rapidchain, cycledger, li2025sp] tolerate a larger fraction (i.e., <1/2<1/2) of malicious nodes within each shard based on a less practical assumption that the network within each shard is synchronous. However, real-world network environments are often not synchronous (e.g., partial-synchronous), especially for large-scale blockchain systems. SGX-Sharding [sgxSharding] reduces the shard size by applying trusted hardware. However, it brings additional deployment overhead to each node, limiting its generality. Some other solutions reduce the shard size at the cost of reducing the systems’ overall resiliency, compromising the system’s security. For example, in Pyramid [pyramid], the assumption entails a smaller fraction of malicious nodes in the network, resulting in a system capable of withstanding less than $1/8$ of malicious nodes.
@@ II-C Blockchain Sharding with Corrupted Shards
Only a handful of blockchain sharding studies allow corrupted shards, with various limitations. Free2Shard [rana2022free2shard] allows corrupted shards and preserves the system’s security via a network-wide consensus. However, it is based on the assumption of a global synchronous network environment. Therefore, it does not work correctly under a practical partial-synchronous network. In addition, Free2Shard adopts the Proof of Work (PoW) consensus protocol, which does not provide deterministic finality and is prone to forking problems. PolyShard [li2020polyshard] leverages coded sharding to enable nodes to store and verify coded shard data, thereby maintaining system security even in the presence of corrupted shards. However, PolyShard also requires an ideal synchronous network environment, which may limit its practicality. Additionally, the computational overhead of polynomial encoding/decoding and the communication costs of network-wide data exchange could impact both its performance and generality. In GearBox [gearbox], the system relies on a network-wide consensus to ensure the system’s security, which can easily lead to a performance bottleneck in large-scale networks. The latest work, known as CoChain [cochain], similarly stems from an intuition that allows shard corruption to enhance concurrency. In CoChain, shards are randomly assigned to CoC (Consensus on Consensus) groups. However, CoChain only permits the failure of less than one-third of the shards within a group. When the group size is no larger than 3, CoChain cannot allow any corrupted shard, losing feasibility. More importantly, CoChain needs to rely on very time-consuming state migration to replace the corrupted shards when they fork. In contrast, our SpiralShard can utilize the LCE protocol to easily help corrupted shards remove excess forks without the need for shard replacement or recovery.
@@
SpiralShard is a novel design that is different from existing blockchain sharding systems. Built upon the assumption of partial-synchronous networks, it is more practical. Leveraging BFT-typed intra-shard consensus ensures finality and energy efficiency. Unlike systems reliant on global consensus to resolve forks, it leverages multiple endorsement groups, preserving scalability and low overhead. Through a series of deliberate design choices–including the LCE protocol and the controlled tolerance of statistically bounded corrupted shards–SpiralShard raises transaction-level concurrency while rigorously preserving system security. Table I provides a qualitative comparison between SpiralShard and several representative blockchain sharding schemes.
TABLE I: Qualitative comparison of representative blockchain sharding protocols.
Protocol Corrupted Shards Network Model Consensus Type Network-wide Protocol Overhead Concurrency
Communication
[elastico, omniledger, pyramid, li2022jenga], etc. Not Allowed Partial-synchronous BFT-type Needless Low Low
[rapidchain, cycledger, repchain, li2025sp], etc. Not Allowed Synchronous BFT-type Needless Low Medium
Free2Shard [rana2022free2shard], PolyShard [li2020polyshard] Allowed Synchronous PoW Need High Medium
GearBox [gearbox] Allowed Partial-synchronous BFT-type Need High Medium
SpiralShard Allowed Partial-synchronous BFT-type Needless Low High
III Network and Threat Model
@@ III-A Network Model
SpiralShard works on a partial-synchronous Peer to Peer (P2P) network, in which there is a known bound $\Delta$ and an unknown Global Stabilization Time (GST). After GST, the network becomes synchronous, and all transmissions between two honest nodes arrive within time $\Delta$. Like most previous blockchain sharding systems, the messages are propagated through gossip protocol. Each node in the system has its public/secret key pair, which can be authenticated through a trusted Public-Key Infrastructure (PKI) to represent its identity while sending messages. Like many prior works [pyramid, omniledger, Harmony], SpiralShard utilize a beacon chain to publicly store the identities of participants.
@@
In SpiralShard, the network consists of $N$ nodes. The shard size (i.e., the number of nodes per shard) and the endorsement group size (i.e., the number of shards per endorsement group) are denoted as $S$ and $G$, respectively. This implies that each endorsement group contains $S\cdot G$ nodes, with $\lfloor N/S\rfloor$ shards within the network. A node’s endorsement group identity is determined by its shard identity (i.e., $GroupID=\lfloor ShardID/G\rfloor$).
@@
SpiralShard employs the account/balance model, widely adopted by existing works to represent ledger states [pyramid, Harmony, monoxide, LBCHAIN, sgxSharding, sharper, brokerchain]. In SpiralShard, a specific account’s state (e.g., balance) is maintained by one shard, determined by the hash of its account address. Accordingly, a transaction in the network is routed to the corresponding shard based on its associated account address. It is important to emphasize that this includes both intra-shard and cross-shard transactions. Besides, the accounts involved in cross-shard transactions may be in different endorsement groups.
III-B Threat Model